OpenStudy (18jonea):

If you lift a bag of flour from the floor to a shelf 1.75 meters high in 3 seconds using 50 W of power, how much force are you applying to the bag? 85.7 N 262.5 N 29.2 N 57.1 N @loser66

OpenStudy (18jonea):

@.Sam.

OpenStudy (18jonea):

@osprey

OpenStudy (osprey):

without wading through the arithmetic on this one (later, perhaps if asked again) I'd go for the lowest number of Newtons. Reason is that 29 N is quite a big force, and I'm thinking of lifting and doing yourself an injury by doing so. Which makes this a GUESSTIMATE.

OpenStudy (18jonea):

i need to be sure on this question could you please help me work it out through numbers @osprey

OpenStudy (18jonea):

@Elsa213

OpenStudy (18jonea):

@mathmate

OpenStudy (mathmate):

Try not to use the answer options. Do your own calculations and then find a match. hint: Power=F*D/t so F=Power*t/D (in suitable units)

OpenStudy (18jonea):

29.16

OpenStudy (osprey):

That's the most realistic one if the calc is right, which I hope it is. @mathmate medal for the 'rith ... hope we're right ...

OpenStudy (18jonea):

it was 85.7

OpenStudy (mathmate):

@18jonea F=power*time/height =50 watts*3 seconds / 1.75 m note 1 watt = 1 N-m/s =50 N-m/s * 3 s / 1.75 m = 50*3/1.75 \(\frac{N-m \times s}{s \times m}\) = 85.7 N so the last answer is correct.